OpenStudy (anonymous):

∣ ¾ x – 2 ∣ - 3 = - 1

OpenStudy (anonymous):

[x = 16/3, _X = 2], [x = 0, _X = 2]

OpenStudy (anonymous):

can you explain how you got 16/3? i got the rest of the equation but don't see where i missed the 16/3 part...maybe i'm having a blonde moment. thank you

OpenStudy (anonymous):

∣ ¾ x – 2 ∣ - 3 = - 1 ¾ x – 2 = 2 and ¾ x – 2 = -2

OpenStudy (anonymous):

oh i see now, i forgot to isolate the absolut value, thanks alot.

OpenStudy (anonymous):

De nada but might wanna dbl check